# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ 2.3 Evaluation metrics and limitations
We now outline the conventional evaluation metrics in survival analysis. The detailed calculation of these metrics is outlined in Appendix E and here we mainly focus on why those metrics are problematic under dependent censoring.
Concordance index (CI) evaluates the ranking of individuals when their relative ordering is of primary interest. It measures the proportion of concordant pairs – pairs where the predicted ranking of risk scores aligns with the true ordering of event times – among all comparable pairs [Harrell Jr et al., 1996]. A pair $(i,j)$ is defined as comparable if $\delta_{i}=1$ and $t_{i}<t_{j}$, which means that the earlier individual has experienced the event and the observed time of the later individual is greater than the first. Even under the random censoring assumption, it is evident that the comparable pairs selected using the above criteria do not represent all possible pairs based on their true event times:
@@
$\begin{array}[]{l}\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})\\ =\quad\text{Pr}(\,e_{i}<c_{i},\ \ e_{i}\ <\ \min\{c_{j},e_{j}\}\ )\\ =\quad\text{Pr}(\,e_{i}<c_{i},\ t_{e,i}<c_{j},\ e_{i}<e_{j})\\ =\quad\underline{\text{Pr}(\,c_{i}<c_{i})\ \text{Pr}(e_{i}<c_{j})}\quad\text{Pr}(e_{i}<e_{j})\ ,\end{array}$
where the factorization follows from the independent censoring assumption. Compared to the right-hand side term, this expression includes additional underlined terms related to the censoring distribution, confirming the findings of Uno et al. [2011].
To address this limitation, Uno et al. [2011] introduced a modified version of CI that employs inverse probability of censoring weights (IPCW), ensuring consistency with the true CI. The IPCW approach estimates the probability of being uncensored at each time point and reweights the uncensored instances accordingly.
Integrated Brier Score (IBS) evaluates the overall accuracy of predicted survival functions by integrating the Brier Score (BS) [Brier, 1950] over a range of time points. It measures the average squared difference between the predicted survival probabilities and actual survival outcomes over time [Gerds and Schumacher, 2006]. To handle censored instances, IPCW can be used to reweight uncensored instances when computing both BS and IBS [Graf et al., 1999]. Similar to IPCW-CI, IPCW weights are often approximated using the KM estimator, which is biased under dependent censoring [Emura and Chen, 2018, Ch. 1].
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [2020] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $c_{i}$, its margin time is calculated as:
@@
where $\hat{S}(t)$ is an estimator, e.g., the KM, derived from the training dataset.
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption and its computational simplicity – requiring no interaction with the covariates – it becomes problematic when censoring and event times are dependent [Emura and Chen, 2018, Ch. 1]. In such cases, the KM estimates may be biased and inconsistent [Campigotto and Weller, 2014]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [Gerds and Schumacher, 2006]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.
@@ 3 The Proposed Metrics
CI-Dep: We propose a modified version of Harrell’s CI [Harrell Jr et al., 1982] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\theta$. When $\theta\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in Uno et al. [2011]. Our proposed concordance index thus requires a known Archimedean copula, that models the dependency between marginals $E$ and $C$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
$\small\begin{split}\text{CI}(\mathcal{D};\,\hat{\eta}(\cdot))=\frac{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.\end{split}$ (3.1)
Now, employing the IPCW technique used by Uno et al. [2011], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(C>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We can then reweight Equation 3.1 using the CG estimator as:
$\small\begin{split}&\text{CI}_{\text{Dep}}(\mathcal{D};\,\hat{\eta}(\cdot))\ =\\ &\frac{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},\end{split}$ (3.2)
@@
IBS-Dep: We propose a modified version of the Brier Score [Graf et al., 1999] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$e^{\text{margin}}(t_{i})=e^{\text{margin}}(\hat{S}_{\text{CG}},t_{i})$, where $\hat{S}_{\text{CG}}$ is the marginal survival probability estimated using the CG estimator. We then formulate the dependent Brier Score as:
$\begin{array}[]{l}\text{BS}_{\text{Dep}}\left(\mathcal{D};\,\hat{S}(\cdot),t^{*}\right)\\ =\frac{1}{N}\sum_{i=1}^{N}\Bigg{(}\mathbbm{1}[\tilde{e}_{i}>t^{*}]-\hat{S}(t^{*}\mid\bm{x}_{i})\Bigg{)}^{2}.\end{array}$ (3.3)
$\text{where}\quad\tilde{e}_{i}=\begin{cases}e_{i}&\text{if}\quad\delta_{i}=1\\ e_{i}^{\text{margin}}(t_{i})&\text{elsewhere.}\end{cases}$
@@
where $t_{\text{max}}$ is usually the maximum observed time.
MAE-Dep: We propose a modified version of the MAE-Margin [Haider et al., 2020] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess”-estimates. Furthermore, for each censored instance, we adopt the approach from Haider et al. [2020] and assign a confidence weight $\omega_{i}=1-\hat{S}_{\text{\text{CG}}}(t_{i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot e^{\text{margin (CG)}}(t_{i})$
$\qquad+\quad\delta_{i}\cdot t_{i}\big{]}-\hat{t}_{i}\Big{|}\ .$ (3.5)
4 Experiments and results
For our empirical analyses, we evaluate the effectiveness of the proposed metrics under dependent censoring in synthetic datasets under a known copula and in semi-synthetic datasets by learning the copula based on the data.
@@ 4.1 Experimental Setup
Figure 2: Outline of the proposed evaluation pipeline. Based on some raw, survival dataset $\mathcal{D}$, we (1) form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. We then (2) select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), and (3) generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. We then (4) create $\mathcal{D}^{\prime\prime}$ based on the ground-truth dataset, the learned censoring distribution, and the selected features. Using $\mathcal{D}^{\prime\prime}$, we (5) train two copula models, i.e., Clayton and Frank, and use AIC to select the copula $C_{\theta}$ that best suits the data. At the same time, we (6) train 5 survival learners independent of the copula on $\mathcal{D}^{\prime\prime}$ to produce survival predictions. Lastly, we calculate the (censored) evaluation error and (7) compare this to the true error.
Synthetic datasets. Following the approach of Foomani et al. [2023], we create a synthetic survival dataset from a data generation process (DGP). The event and censoring distributions are specified by linear Weibull marginal distributions. Specifically, we sample R.V.s $E$ and $C$ from a known copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula, e.g., Clayton or Frank, known at evaluation time. See Appendix F.1 for more details.
Semi-synthetic datasets. To evaluate our proposed metrics, we need to know the true metric, which requires precise knowledge of when each instance experiences the event. Due to censoring, event times are not available in real-world datasets for all instances, so we create a realistic, semi-synthetic dataset, where the covariates, event distribution, and censoring distribution align with some real-world datasets. This approach is inspired by Qi et al. [2023a], who proposed a semi-synthetic framework to generate different kinds of censoring distributions. Figure 2 shows the pipeline to the create semi-synthetic datasets.
In summary, based on some real-world survival dataset $\mathcal{D}$, we calculate the censoring distribution $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$, create $\mathcal{D^{\prime}}$ by removing the censored instances from $\mathcal{D}$, and then apply the censoring distribution to $\mathcal{D^{\prime}}$, so that for each instance, that instance is synthetically censored if the censoring time is earlier than the event time ($\tilde{c}_{i}<t_{i}$), and otherwise is left uncensored. We perform a feature selection step on $\mathcal{D}$ to induce dependent censoring using one of three strategies:
  • Original, where the original feature set is kept intact.
  • Top-$k$, where we select the $k$ most important features from $\mathcal{D}$ based on their predictive power. We train a CoxPH model on $\mathcal{D}$ and then use permutation feature importances [Breiman, 2001] to select the $k$ best features. Feature importances are calculated by shuffling feature values and observing how much the model’s performance degrades, using Harrell’s CI.
  • Random-%, where we select a random proportion of %-pct features from the original feature set. In a small dataset, randomly selecting 25% of the features can introduce significant residual dependence if key features are excluded. In contrast, in a high-dimensional dataset, this risk is lower as many important features are likely to remain.
Finally, we have $\mathcal{D}^{\prime\prime}$: a dataset with a reasonable, but synthetic censoring distribution with a restricted feature set and where the ground-truth event times are known. This will allow us to compare the proposed and existing (censored) metrics with the ground-truth metrics.
Copula model. To validate our dependent metrics in semi-synthetic datasets, we train two copula models, Clayton and Frank, and two nonlinear Weibull proportional-hazards models for the marginal distributions, $E$ and $C$, respectively, on $\mathcal{D}^{\prime\prime}$. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture with a single hidden layer and the nonlinear ReLU activation function. For each copula model, we train the marginals and the copula $C_{\theta}$ jointly using maximum likelihood estimation (MLE) (see Appendix F.3 for details). After training the two candidate models, we choose the model that minimizes the Akaike Information Criterion (AIC), given its loss on the validation set and number of parameters. This offers a practical, data-driven approach to copula selection [Emura and Chen, 2018, Ch. 4] by selecting the copula model with the lowest validation loss. We then use the best copula $C_{\theta}$ for the evaluation of the dependent metrics.
Datasets. We apply the evaluation pipeline in Figure 2 to 6 large, real-world datasets: METABRIC [Curtis et al., 2012], 2 from MIMIC-IV (all causes and hospital death) [Johnson et al., 2023], and 3 from SEER (death from brain, liver and stomach cancer) [LAG et al., 2007]. They differ in their number of instances, features, and censoring rates. Appendix F.1 gives further details on the datasets and describes how we preprocessed them.
Survival learners. We compare the prediction results using 5 survival learners produced by the CoxPH [Cox, 1972], gradient boosting (GBSA) [Ridgeway, 1999], Random Survival Forest (RSF) Ishwaran et al. [2008], DeepSurv Katzman et al. [2018] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. These learners predict individual survival distributions (ISDs) for each instance $i$. See Appendix F.2 for further details, including implementation details and hyperparameters.
We train the five survival learners across 5 experiments given a random seed (0-4). In each experiment, we follow Sechidis et al. [2011] and first split the dataset into stratified training (70%), validation (10%), and test (20%) sets. After splitting, we impute missing values using sample mean for real-valued covariates or mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy. All results are reported in the test sets.
Figure 3: Plot of the bias ($Y$-axis) – i.e., the proximity to the true error – of established baseline metrics and the proposed dependent metrics as a function of the dependency ($X$-axis) under a known copula. We plot the mean biases ($\pm$ SD.) in the test sets across 100 experiments in a synthetic dataset ($N=10,000$, $d=10$). Overall, our dependent metrics show lower biases when dependency is notable ($\tau>0.25$) than the independent metrics. See Appendix F.1 for details.

### v3

@@ 2.3 Evaluation metrics and limitations
We now outline the conventional evaluation metrics in survival analysis and discuss why these metrics are problematic under dependent censoring. See Appendix D for their formal definitions.
Concordance index (CI) evaluates the ranking of individuals when their relative ordering is of primary interest. It measures the proportion of concordant pairs – pairs where the predicted ranking of risk scores aligns with the true ordering of event times – among all comparable pairs [29]. A pair $(i,j)$ is defined as comparable if $\delta_{i}=1$ and $t_{i}<t_{j}$, which means that the earlier individual has experienced the event and the observed time of the later individual is greater than the first. Even under the random censoring assumption, it is evident that the comparable pairs selected using the above criteria do not represent all possible pairs based on their true event times:
@@
$\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})$ $=\quad\text{Pr}(\,e_{i}<c_{i},\ \ e_{i}\ <\ \min\{c_{j},e_{j}\}\ )$
$=\quad\text{Pr}(\,e_{i}<c_{i},\ t_{e,i}<c_{j},\ e_{i}<e_{j})$
$=\quad\underline{\text{Pr}(\,c_{i}<c_{i})\ \text{Pr}(e_{i}<c_{j})}\quad\text{Pr}(e_{i}<e_{j})\ ,$
where the factorization follows from the independent censoring assumption. Compared to the right-hand side term, this expression includes additional underlined terms related to the censoring distribution, confirming the findings of [30]. To address this limitation, Uno et al. [30] introduced a modified version of CI that employs inverse probability of censoring weights (IPCW), ensuring consistency with the true CI. The IPCW approach estimates the probability of being uncensored at each time point and reweighs the uncensored instances accordingly.
Integrated Brier Score (IBS) evaluates the overall accuracy of the predicted survival functions by integrating the Brier Score (BS) [21] over a range of time points. It measures the average squared difference between the predicted survival probabilities and actual survival outcomes over time [31]. To handle censored instances, IPCW can be used to reweigh the uncensored instances when computing both BS and IBS [32]. Similar to CI-IPCW [30], IPCW weights are often approximated using the KM estimator, which is biased under dependent censoring [4, Ch. 1].
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [22] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $t_{i}$, using some estimator $\hat{S}(t)$, e.g., the KM, the subject’s margin time is calculated as:
@@
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption, it becomes problematic when censoring and event times are dependent [4, Ch. 1]. In such scenarios, the KM estimates may be biased and inconsistent [33]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [31]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.
@@ 3 The proposed metrics
CI-Dep: We propose a modified version of Harrell’s CI [20] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\hat{\theta}$. When $\hat{\theta}\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in [30]. Our proposed concordance index thus requires a known or approximated Archimedean copula that models the dependency between marginals $T$ and $C$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
$\begin{split}\text{CI}(\mathcal{D};\,\hat{\eta}(\cdot))=\frac{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.\end{split}$ (3.1)
Now, employing the IPCW technique used by [30], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(C>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We then reweigh Equation 3.1 using the CG estimator as:
$\begin{split}\text{CI}_{\text{Dep}}(\mathcal{D};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},\end{split}$ (3.2)
@@
IBS-Dep: We propose a modified version of the Brier Score [32] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given that the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$e^{\text{margin}}(t_{i})=e^{\text{margin}}(\hat{S}_{\text{CG}},t_{i})$, where $\hat{S}_{\text{CG}}$ is the marginal survival probability estimated using the CG estimator, then we have:
$\begin{array}[]{l}\text{BS}_{\text{Dep}}\left(\mathcal{D};\,\hat{S}(\cdot),t^{*}\right)=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbbm{1}[\tilde{e}_{i}>t^{*}]-\hat{S}(t^{*}\mid\bm{x}_{i})\right)^{2},\end{array}$ (3.3)
$\text{where}\quad\tilde{e}_{i}=\begin{cases}e_{i}&\text{if}\quad\delta_{i}=1\\ e_{i}^{\text{margin (CG)}}(t_{i})&\text{elsewhere.}\end{cases}$
@@
where $t_{\text{max}}$ is usually the maximum observed time in the training set.
MAE-Dep: We propose a modified version of the MAE-Margin [22] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess” estimates. Furthermore, for each censored instance, we adopt the approach from [22] and assign a confidence weight $\omega_{i}=1-\hat{S}_{\text{\text{CG}}}(t_{i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot e^{\text{margin (CG)}}(t_{i})+\delta_{i}\cdot t_{i}\big{]}-\hat{t}_{i}\Big{|}.$ (3.5)
4 Synthetic experiments
@@ 4.1 Experimental setup
We create synthetic datasets from a data generation process (DGP) under the Weibull distribution (see Appendix E.1). We sample R.V.s $T$ and $C$ from a copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula. In each of 100 experiments, we follow [34] and split the data into stratified training (70%), validation (10%), and test (20%) sets (seeds 0-99). Here, we investigate our metrics under dependency and censoring. See Appendix F.1 for evaluating under copula misspecification and varying sample sizes.